three blocks m1,m2 and m3 of masses 8kg,3kg and1 kg are placed in contact on a  smooth surface.Forces F1=140N and F2=20N  are acting on blocks m1 and m3 respectivly.The reaction between blocks m2 and m3 is
1)2.5N
2)7.5N
3)22.5N
4)30N
 

Grade:12th pass

1 Answers

Pooja
4 years ago
                  T2       T1
20N⬅️1⃣➖3⃣➖8⃣➡️140N
             m3     m2    m1
                          ➖➖➡️a  
The net force on the system is 140-          20=120N
So acceleration is 120÷(8+3+1)=10m/s2
Let the force between m3 and  m2 be  T2
 Using free body diagram of m3
                    ➖➡️m3(a)
          20⬅️1⃣➡️T2
 T2-20=m3(a)
Hence T2= 1(10)+20
          T2 = 30N
The reaction between blocks m2 and m3 is 30 N (4th option)